AGw Commissioning with IRTC and TCS Software Testing - 30-31 October MST

Observers: J. Hill on mtn, A. Rakich in Tucson, N. Ageorges in Garching
Software Support: N. Cushing in Tucson plus T. Leibold and J. Borelli at beginning
Telescope Operator: D. Gonzalez Huerta

Summary

We tested the new PSF scheme to automatically offload active optics corrections from M2 to M1. Images are good after the offload. And it executes as advertised when you issue a stepFocus. It seems to be ready for production for LUCIFER observing. So the wavefront corrections should be sent to PSF with "default" setting on the PSF Wavefront sub-gui.

We tested (and used) the new "ACQUIRE" movie mode in GCS - sweet. We didn't do all that many guiding presets, but GCS didn't miss any stars that landed on the acquisition frame.

We had a wild problem with both primary mirrors oscillating by millimeters. It seems to be stimulated by the PMC Elevation Feed Forward being switched on and off as fast as several times per second. We also learned that the "buzzing" in the left image seen in previous weeks is caused by the right primary mirror oscillating up and down.

We worked some more on measuring the IRTC focus, but the activity was fairly well destroyed by the primary mirror oscillations.

We took 36 IRTC images to make a pointing model for the last three hours of the night.

Details

Times reported in UT below

There were no local capillaries on the E1 elevation pad tonight.

Open

00:55 Open Dome. Sky mostly clear with clouds to SW. Wind light. T=6.7 degC D=-2 degC.

01:15 update focus in AIP_L.cfg from 29 to 32 to make Doug's estimated focus adjustment. (Last night's late changes didn't make it into today's TCS build.) Active optics gain is back to the default 0.8.

01:20 David is reconnecting to the AGw software

01:23 IE=+5 CA=-38 to start where we ended last night.

01:24 pointto M5_1568 gs=6 TelM="ACQUIRE" PAM- (don't send a guide star in parallactic mode)

01:26 pointto M5_1568 TelM="TRACK" PAM- -- a comatic star appears at the top of IRTC Apply Z=+0.2.

01:30 Adjust IE=+17 CA=-37

Testing new parts of TCS software, and generally collimating around

01:32 pointto M5_1568 gs=6 TelM="ACTIVE"

The first active optics correction to "default" (which put focus and coma on M2) made the image worse. I switched Wavefront correction to "M1", and then it seems to converge acceptably in 5 iterations.

01:42 Set Wavefront correction to "default". Ran a few corrections. Then sent a "focus 0" from IRAF. As advertised, this removed the "Active Optics" vector from the M2, and applied the equivalent correction to the "Secondary Offload" vector on M1. The image had the usual arcsec of wobbling when the primary makes a significant move, but otherwise seemed unchanged. The next active correction from GCS was modest, so it seems like everything worked OK.

Seeing is around 1 arcsec, and active corrections are bouncing around at the 500-700 nm level.

02:01 Remove active corrections from M1. Allow corrections to reappear on M2 with active optics.

02:06 focus 0 again.

02:09 Jose is starting a multi-client test with irs.

02:11 pointto M5_1568 gs=2 TelM="ACTIVE" - The preset succeeded with a warning from GCS that the guide star was too faint. (The in-focus star is 4000 counts peak above background, so I'm not sure about this warning.)

LBTtools.IRTC.pointto reported:
   irc PresetTelescope 224.25 POSITION ACTIVE left
   > Problem with PresetTelescope command.  res=2

...Done! (failed!)
   to examine the log: page LOGs/pointto_021110.log

It seems that Dave is incorrectly interpreting the result "2" as a failure when it was only a warning. Jose says 0=success 1=failed 2=warning.

GCSL reported back to irs:
Oct 30 19:12:14 lbtmu05 irc[7679]: couldn't adjust exposure time to meet low counts requirement. star too dim or guide rate too high.
Oct 30 19:12:14 lbtmu05 irc[7679]: maximum exposure time is 4600 ms at the given guide rate, estimated counts are 971
Oct 30 19:12:14 lbtmu05 irc[7679]: adjusted exposure time to 4600 ms

02:27 Guider image is distinctly elongated. Spots on left side of wfsc 0051 pupil are distinctly elongated while spots on right side of pupil are round. What does that mean?

02:29 Adjust M1 Zglobal to +0.3 (to tweak the system with some focus). That did make the left side spots look better.

02:33 Clear C00 on M1 to get rid of the higher order corrections. This didn't really change very much. The guide star still looks elongated after a few iterations. See also the elongated spots in wfsc 0060.

02:39 pointto M5_1568 gs=3 TelM="ACTIVE" (failed - apparently Jose had killed my authorized IIF object during his tests)

Oct 30 19:39:29 lbtmu05 irc[7913]: PresetTelescope execution failed: Requested side not allowed for this instrument / focal station combination

02:40 authorize IRTC

02:42 pointto M5_1568 gs=3 TelM="ACTIVE" - succeeded. The guide star looks much better now, so perhaps g* 2 was a binary. Guider images now improved to 0.5-0.6 arcsec FWHM.

02:50 focus 0

02:55 pointto M5_1568 gs=2 TelM="ACTIVE" -

Now gs 2 seems nice and round. So it isn't a binary. Must have been an active correction problem of some kind.

Collimating with IDL

03:00 stopWFS

03:08 agw_collimate, EXP=30 (fails because I was LBTO@lbtmu05 which doesn't have TCS)

IDL> agw_collimate, EXP=30
Acquiring WFSC Image: should take about 30 seconds....
readWFSCam -e 30000
Error finding new file name
    Error: agw_wfsc_image failed

The error message is obscure, but basically correct.
Oct 30 20:08:53 lbtmu05 LBT: Command called on inactive subsystem GCSL
Oct 30 20:08:53 lbtmu05 LBT: cmd error handling...
Oct 30 20:08:53 lbtmu05 LBT: Command called on inactive subsystem GCSL
Oct 30 20:08:53 lbtmu05 LBT: cmd error handling...

03:12 Try readWFSCam directly:
[telescope@to-station 081031]$ readWFSCam -e 30000
readWFSCam: acquiring full wfsc image ...
GCS side: L
selected AGW: AIP_LUCIFER
exposureTime = 30000 ms
new image (1/1) received: /Repository/IRTC_Data/081031/wfsc000001.fits

03:23 agw_collimate, EXP=30 from lbtmu107

03:25 agw_collimate_loop, EXP=30

% CHECK_FITS: NAXIS keywords in FITS header have been updated
    ...Done: Filename =  /Repository/IRTC_Data/081031/wfsc000004_pupil.fits
=> Processing file wfsc000004_dd.fits
       0       388.43823      247.424
% Type conversion error: Unable to convert given STRING to Float.
% Detected at: ZERNIKE_ROTATION_SETUP   30
  /home/LBTO/agw/idl/zernike/zernike_rotation_setup.pro
       Zern_Rot_Del =   -25.68
            WFSC        Rotated
Zern #   Orientation   to Primary
---------------------------------
Z[ 2] =    -462.70      -247.32
Z[ 3] =     391.51       553.36
......................................................................................etc.
Z[22] =      -9.41        -9.41
% Variable is undefined: FILENAME.
% Execution halted at: ZERNIKE_ROTATION_SETUP   40
  /home/LBTO/agw/idl/zernike/zernike_rotation_setup.pro
%                      AGW_COLLIMATE_LOOP  108
  /home/LBTO/agw/idl/agw/agw_collimate_loop.pro
%                      ZERNIKE_ROTATION_SETUP   40
  /home/LBTO/agw/idl/zernike/zernike_rotation_setup.pro
%                      AGW_COLLIMATE_LOOP  108
  /home/LBTO/agw/idl/agw/agw_collimate_loop.pro
%                      $MAIN$

Apparently the call to zernike_rotation_setup.pro had changed, and had not been updated in agw_collimate_loop.pro. I believe that I have fixed it. Doug: Please check that I used the correct filename variable as they were different in agw_collimate versus agw_collimate_loop.

03:36 agw_collimate, EXP=30 wfsc 0005 (I haven't updated collimation in a while, so the first correction is bogus relative to the collimation table.)

04:00 agw_collimate_loop, EXP=30 (now working with updated code)

04:08 focus 0

04:08 agw_collimate_loop failed while retrieving parameters

Z[11] =     132.79
/Repository/IRTC_Data/081031/wfsc000017_irc.zern
Acquiring WFSC Image: should take about 30 seconds....
readWFSCam -e 30000
    ...Done: Filename =  /Repository/IRTC_Data/081031/wfsc000018.fits
Acquiring Data Dictionary Information...
   Error retrieving DD values
    Stopping collimation routine after irc_getparameter

04:15 focus 0

04:17 agw_collimate_loop, EXP=30

04:27 stopped agw_collimate_loop

ACQUIRE mode tests

04:30 pointto M5_1568 gs=2 TelM="ACQUIRE" - succeeded
GCSGUI brought up a pop-up window with the acquire movie mode running - slick.
Acquire mode does not seem to be saving images to disk.

04:37 Updated IE=+15 CA=-30. It is interesting that IE seems to be getting monotonically more positive over the last few days. What could cause this? Is this a temperature effect -- Mirror Cell temp is now 7.08 degC.

04:39 pointto M5_1568 gs=4 TelM="GUIDE"

Note that different guide stars on the WFS hotspot are landing at different locations on the IRTC with a spread of +-4 arcsec or so. I think this means that the AGw calibrations scheduled for the weekend are important.

04:45 startWFS to start GCS WFSing.

04:46 GCSL is having trouble getting the WFS image, and thus the collimation is falling apart.

Oct 30 22:03:52 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 37716 ms.
Oct 30 22:03:56 lbtmu107 LBT_left_GCS: PROFILING: WFS: triggerWFS() entered.
Oct 30 22:03:56 lbtmu107 LBT_left_GCS: PROFILING: WFS: requesting image...
Oct 30 22:03:56 lbtmu107 LBT_left_GCS: WFS: requesting a WFS image...
Oct 30 22:03:56 lbtmu107 LBT_left_GCS: WFS: intending to wait 37716 ms for the WFS image.
Oct 30 22:05:18 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 37716 ms.
Oct 30 22:05:22 lbtmu107 LBT_left_GCS: PROFILING: WFS: triggerWFS() entered.
Oct 30 22:05:22 lbtmu107 LBT_left_GCS: PROFILING: WFS: requesting image...
Oct 30 22:05:22 lbtmu107 LBT_left_GCS: WFS: requesting a WFS image...
Oct 30 22:05:22 lbtmu107 LBT_left_GCS: WFS: intending to wait 37716 ms for the WFS image.
Oct 30 22:06:44 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 37716 ms.

05:08 stopWFS (GCSL responded correctly to this command)

05:09 pointto M5_1568 gs=4 TelM="ACTIVE" (preset hung, GCSL doesn't seem to be responding - no syslog activity)

05:14 kill and restart GCSL

05:18 pointto M5_1568 gs=4 TelM="ACTIVE" - succeeded

05:24 pauseGuiding

05:26 testing an old IRAF script inex.cl "cl < inex.cl". It updated the primary mirror collimation correctly, but failed updating the M2 position:
irc Move 0. 0. 0. 0. 0. 0. 63 ABS M2 0 left
This reported success to IRAF, but didn't seem to do anything to M2. The socket errors are caused by the IRTC still being in video (grab) mode, but the thing that confuses me is how this error can occur before the previous command in my script completed.

From the irc syslog on lbtmu05:
Oct 30 22:26:08 lbtmu05 irc[9881]: GetFOV result status: OK Field of view (FOV): Wide Fov
Oct 30 22:26:08 lbtmu05 irc[9904]: SetIRParams command result: ERROR
Oct 30 22:26:08 lbtmu05 irc[9904]: [Error] IRSocket::SetIRParams -> IR System busy. Command has been understood, but NOT executed.
Oct 30 22:26:24 lbtmu05 irc[9959]: Move X=0. Y=0. Z=-0.6 RX=0. RY=0. RZ=0. DFLAG=4 MOVETYPE=ABS OPE=M2 TIME=0 Side=left
Oct 30 22:26:24 lbtmu05 irc[9959]: Move result status: OK
Oct 30 22:26:24 lbtmu05 irc[9959]: Ok
Oct 30 22:26:24 lbtmu05 irc[9959]: left PSF Move complete
Oct 30 22:26:27 lbtmu05 irc[9998]: GetFilter result status: OK Filter: H Wavelength: 1.6 [microns]
Oct 30 22:26:27 lbtmu05 irc[10021]: GetFOV result status: OK Field of view (FOV): Wide Fov
Oct 30 22:26:28 lbtmu05 irc[10044]: SetIRParams command result: ERROR
Oct 30 22:26:28 lbtmu05 irc[10044]: [Error] IRSocket::SetIRParams -> IR System busy. Command has been understood, but NOT executed.
Oct 30 22:26:42 lbtmu05 irc[10099]: Move X=0. Y=0. Z=0. RX=0. RY=0. RZ=0. DFLAG=63 MOVETYPE=ABS OPE=M2 TIME=0 Side=left
Oct 30 22:26:42 lbtmu05 irc[10099]: Move result status: OK
Oct 30 22:26:42 lbtmu05 irc[10099]: Ok
Oct 30 22:26:42 lbtmu05 irc[10099]: left PSF Move complete

My script:
# J. M. Hill, LBT Observatory
# 12 April 2008

# Assume that the telescope is already tracking on the star

# update the M1 collimation
moveOPE M1 X=0.0 Y=0.0 Z=0.0 RX=0.0 RY=0.0 RZ=0.0 ABS+ Magic-

# move to the intra-focal position
moveOPE M2 X=0.0 Y=0.0 Z=0.6 RX=0.0 RY=0.0 RZ=0.0 ABS+ Magic-

# take a picture
takepic exptime=1.5

# move to the extrafocal position
moveOPE M2 X=0.0 Y=0.0 Z=-0.6 RX=0.0 RY=0.0 RZ=0.0 ABS+ Magic-

# take another picture
takepic exptime=1.5

# put the secondary back for good housekeeping (must use Magic+)
moveOPE M2 X=0.0 Y=0.0 Z=0.0 RX=0.0 RY=0.0 RZ=0.0 ABS+ Magic+

05:40 take a test cube with LBTtools.IRTC.takeseq and it worked.

05:53 pauseGuiding (but guiding didn't pause); focus 0

05:54 cl < inex_seq.cl (had some typos in the size of the offsets, and guiding didn't pause, but got the data )
-rw-r--r--    1 root      root         172800 Oct 30 22:54 irtc.20081031.055453.fits
-rw-r--r--    1 root      root         172800 Oct 30 22:55 irtc.20081031.055533.fits
-rw-r--r--    1 root      root         172800 Oct 30 22:56 irtc.20081031.055619.fits

Note: the permissions on the IRTC images have gotten messed up on the new TCS

05:59 pauseGuiding (worked - actually paused)

06:00 resumeGuiding

06:09 pauseGuiding (worked)

The nearly conventional focus sequence

06:28 pointto M5_0094 gs=2 ACTIVE x=0 y=-74.4 focus=29.237

06:37 Stop PMCL, PMCR and ECS to get new versions that will connect to new telemetry. And we are moving PMCR from lbtmu107 to lbtmu105 to lower the load on lbtmu107 as the GCS wfs cycle time is more than 60sec.

06:45 focus 0 to offload corrections from M2

06:51 takeseq 0.006 1000 frmrate=50 take1+ cube 0099

Oct 30 23:51:29 lbtmu05 irc[12146]: GetFilter result status: ERROR [Error] IRSocket::GetFilter::res -> Could not read from socket.

06:52 takeseq 0.006 1000 frmrate=50 take1+ (this worked, but it had an active optics update in the middle of the cube)

06:54 pauseGuiding

06:55 takeseq 0.006 1000 frmrate=50 take1+ cube 0101

06:59 resumeGuiding, setFocus 32.237 (equivalent to on-axis 35)

07:06 pauseGuiding

07:07 takeseq 0.006 1000 frmrate=50 take1+

Oct 31 00:07:11 lbtmu05 irc[12522]: GetFilter result status: ERROR [Error] IRSocket::GetFilter::res -> Could not read from socket.

07:08 takeseq 0.006 1000 frmrate=50 take1+ cube 0101

07:11 resumeGuiding, setFocus 35.237 (equivalent to on-axis 38)

Primary mirror is oscillating by millimeters. See also wfsc 0132 which seems to indicate that this was happening earlier. The problem seems to be that PMCL (and PMCR also) are turning Elevation Feed Forward on and off at rates of 1-10 Hz (it looks random in time). These on/off events seemed to start at 23:51 MST.

07:xx PMCR crashes

07:34 Restart MCSPU

07:36 pointto M5_0094 gs=2 ACTIVE x=0 y=-74.4 (rotator not ready)

07:37 pointto M5_0094 gs=2 ACTIVE x=0 y=-74.4 (no polys)
Oct 31 00:37:32 lbtmu05 irc[13046]: MCSPU error returned - ERROR :  no polys are being received!

07:40 pointto M5_0094 gs=2 ACTIVE x=0 y=-74.4 (no polys)

07:41 David restarts PCS.

07:43 pointto M5_0094 gs=2 ACTIVE x=0 y=-74.4 (didn't find star, we forgot to restore IE, CA)

07:44 pointto M5_0094 gs=2 ACTIVE x=0 y=-74.4 (still we didn't find a star at EL=83)

Move to a star at lower elevation

07:47 pointto M5_0151 gs=2 ACTIVE (didn't find a star here either)

07:51 pointto M5_0151 gs=0 ACQUIRE set IE=48 CA=-30

07:55 pointto M5_0151 gs=2 ACTIVE x=0 y=-113.9 focus=25.610 - success

Feedforward on/off events have resumed on PMCL.

Fri Oct 31 01:02:54.630 2008  PMCR right PMC Stopping elevation angle feed forward
Fri Oct 31 01:02:54.683 2008  PMCR right PMC Starting elevation angle feed forward
Fri Oct 31 01:02:54.775 2008  GCSL [gcs.command.getGuideImage.triggered]
Fri Oct 31 01:02:54.776 2008  PMCL left PMC Stopping elevation angle feed forward
Fri Oct 31 01:02:54.829 2008  PMCL left PMC Starting elevation angle feed forward
Fri Oct 31 01:02:54.991 2008  PMCR right PMC Stopping elevation angle feed forward
Fri Oct 31 01:02:55.028 2008  PMCR right PMC Starting elevation angle feed forward
Fri Oct 31 01:02:55.250 2008  ECS Instrument cooling command set temperature set pt completed:temp 5.8123336e+00 
Fri Oct 31 01:02:56.062 2008  PCS PCSApplyGuideCorrection started - guideX -2.9480568e-01 and guideY 1.1398559e+02
Fri Oct 31 01:02:56.094 2008  PCS PCS ApplyGuideCorrection complete - guideX -2.9480568e-01 and guideY 1.1398559e+02
Fri Oct 31 01:02:56.117 2008  GCSL [gcs.command.getCentroid.triggered]
Fri Oct 31 01:02:56.145 2008  ECS Hydraulic bearing command set temperature set point completed:temp 4.8356667e+00 
Fri Oct 31 01:02:56.189 2008  GCSL [gcs.command.getCentroid.triggered]
Fri Oct 31 01:02:56.271 2008  GCSL [gcs.command.getCentroid.triggered]
Fri Oct 31 01:02:56.655 2008  PMCL left PMC Stopping elevation angle feed forward
Fri Oct 31 01:02:56.707 2008  PMCL left PMC Starting elevation angle feed forward
Fri Oct 31 01:02:56.898 2008  GCSL [gcs.command.getCentroid.triggered]
Fri Oct 31 01:02:57.190 2008  GCSL [gcs.command.getCentroid.triggered]
Fri Oct 31 01:02:57.262 2008  GCSL [gcs.command.getCentroid.triggered]
Fri Oct 31 01:02:57.319 2008  GCSL [gcs.command.getCentroid.triggered]

GCS reported cycle time is up to 64 - 75 sec for the normal 30 sec exposure. lbtmu107 has a lot of swap space in use. Presently GCSL and PSFR are running along with 4 versions of Guider.py and one DDBrowser

08:08 pauseGuiding; takeseq 0.0025 1000 frmrate=50 take1+ 0102

08:11 setFocus 28.61; resumeGuiding

Oct 31 01:16:38 lbtmu05 irc[14261]: GetFilter result status: ERROR [Error] IRSocket::GetFilter::res -> Could not read from socket.

08:16 pauseGuiding; takeseq 0.0025 1000 frmrate=50 take1+ =cube 0103
Image is buzzing back and forth by 2 arcsec p-v in this cube. Is this buzzing caused by the PMC FeedForward Issue.

08:19 Stop loops on PMCL. Image continues to buzz wildly.

08:20 Stop FF on PMCL. Image continues to buzz.

08:24 azhold, elhold FF problems continue on PMCR (David has FF disabled on PMCL)

08:25 resume tracking; disable FF on both mirrors.

08:27 restarting FF on right mirror makes the left image buzz.

08:21 Mirror starts bouncing when David restarts the loops. He gets is under control in a minute or so.

Historical event logs are in /home/ncushing.

08:30 left image does not buzz when left FF is on and right FF is off. What does that mean?

08:32 Stop left primary mirror ventilation, and then right mirror ventilation. Made no difference in image.

08:33 Turn on Right FF - no buzzing

08:34 Collimate Right M1 - no buzzing

08:36 Turned on Mirror Ventilation and the buzzing started. The symptoms are also easily visible on the PMCR hardpoint GUI.

08:37 Lowering the Right primary mirror

08:41 resumeGuiding returned error
resumeGuiding failed!

We think this was because we stopped tracking and the preset state deescalated.

08:42 pointto M5_0151 gs=2 TelM="ACTIVE"

08:43 setFocus 31.61

08:44 Stop and start PMCL FF several times. This makes no perceptible difference to IRTC images being taken at 50 Hz.

08:47 focus 0 for M2 offload

08:48 SX primary oscillates and panics during a collimation update.

08:58 Lowered kP gains to 1.0 for FZ, MX, MY loops on SX, and it seems willing to move in a stable fashion. Or perhaps it is stable because the FF on/off events have slowed to a low rate.

09:00 pointto M5_0151 gs=2 TelM="ACTIVE"

09:03 setFocus 31.61

09:03 PSFL crashes for no obvious reason.

09:06 Stopped one of the 3 GCSGUI that were running. This seems to lower the cycle time from 71 to 64 sec, but that may have been coincidence.

09:10 focus 0

09:14 pauseGuiding; takeseq 0.0025 1000 frmrate=50 take1+ =one cube 0104

09:20 cl < inex_seq.cl 200 x 250 msec at 50 Hz -- three cubes (intra/extra/sky) 0105, 0106, 0107

09:26 setFocus 34.61; resumeGuiding; This image is clearly a donut on IRTC, so we have moved above focus (IRTC is extrafocal). Guider images are 0.6 arcsec.

09:35 pauseGuiding; takeseq 0.0025 1000 frmrate=50 take1+ =one cube 0108

09:39 cl < inex_seq.cl 200 x 250 msec at 50 Hz -- three cubes (intra/extra/sky) but only got 0109 and 0110

09:47 setFocus 26.61; resumeGuiding

09:52 pauseGuiding; takeseq 0.0025 1000 frmrate=50 take1+ =one cube 0111

09:56 cl < inex_seq.cl 200 x 250 msec at 50 Hz -- intra cube is 0112
The hexapod went from +0.6 to -0.8 not the +0.7 to -0.7 like we asked. That was because there was a bit of active optics focus correction on M2.
And it didn't seem to take the extrafocal cube. So there may be a script problem or something here. It seems to be our friend the socket error.
Oct 31 02:58:30 lbtmu05 irc[17643]: GetFilter result status: ERROR [Error] IRSocket::GetFilter::res -> Could not read from socket.

10:04 resumeGuiding, but lost star. SX primary was oscillating again under rapid FF changes.

10:10 Lowered loop gain FZ, MX, My to 0.8, and stopped PMCL elevation FF and guiding restarts when the star reappears.

10:15 pauseGuiding; akeseq 0.0025 1000 frmrate=50 take1+ =one cube 0113

10:19 cl < inex_seq.cl 200 x 250 msec at 50 Hz intra=0114 extra=0115 sky=0116

Collecting Data for Pointing Model with IRTC

10:23 Turn on pointing logging in PCS. Log file is /home/telescope/20081031. Images will be in /Repository/IRTC_Data/081031/.

10:25 focus 0

Starting model terms for data collection
*** Pointing Model Terms ***
IA      -286.93580
IE        48.00000
NPAE      17.99860
CA       -30.00000
AW         5.10060
AN        24.39890
TF        21.06980

10:28 pointto BS9116 TRACK
10:31 takepic 1.5 irtc.20081031.103117.fits

BS9108 irtc.20081031.103504.fits

Oct 31 03:34:24 lbtmu05 irc[19662]: GetFilter result status: ERROR [Error] IRSocket::GetFilter::res -> Could not read from socket.
Oct 31 03:34:32 lbtmu05 irc[19701]: GetFilter result status: ERROR [Error] IRSocket::GetFilter::res -> Could not read from socket.
Oct 31 03:34:44 lbtmu05 irc[19740]: GetFilter result status: ERROR [Error] IRSocket::GetFilter::res -> Could not read from socket.

Blank Sky irtc.20081031.103546.fits

BS9107 irtc.20081031.103831.fits

BS9105 irtc.20081031.104113.fits

BS9104 irtc.20081031.104404.fits (fuzzy star at top edge of picture, EL=10)

BS9107 irtc.20081031.104744.fits

BS9108 irtc.20081031.105026.fits (in the crystal feather)

BS9116 irtc.20081031.105526.fits

BS9118 irtc.20081031.105758.fits

BS9122 irtc.20081031.110214.fits

BS9131 (star not visible in field, and a short spiral search didn't find it) no image

BS9139 irtc.20081031.111003.fits

BS9142 irtc.20081031.111334.fits

BS9145 (star not visible in field, and a short spiral search didn't find it) no image

Eventually had to kill my IRAF session to fix this problem.
IRTC> pointto BS9145 TelM="TRACK"
ERROR: cannot open `uparm$pipe23237' for writing
  "print ('!date -u "+HELPM%S"') | cl | scan (systime)"
     line 153: IRTC$pointto.cl
     called as: `pointto (sourceID=BS9145, TelMode=TRACK)'
IRTC>

BS9148 irtc.20081031.112443.fits (near the tip of the crystal feather)

BS9143 irtc.20081031.112754.fits

11:30 (some of my X-windows quit including IRTC)

BS9141 irtc.20081031.113150.fits

BS9138 irtc.20081031.113346.fits

BS9135 irtc.20081031.113723.fits (in the small crystal feather)

BS9134 irtc.20081031.114049.fits

BS9133 irtc.20081031.114422.fits

BS9126 irtc.20081031.115030.fits (EL=85)

BS9123 irtc.20081031.115413.fits

BS9118 irtc.20081031.115652.fits

BS9116 irtc.20081031.115913.fits

BS9108 irtc.20081031.120222.fits

BS9107 irtc.20081031.120559.fits

BS9122 irtc.20081031.121004.fits

BS9139 irtc.20081031.121358.fits

BS9142 irtc.20081031.121629.fits

BS9148 irtc.20081031.121904.fits

BS9126 irtc.20081031.122440.fits (EL=86.3)

Note: Toggling the IRTC "Grab" mode button on and off is a good way to clear repeated socket errors. Of course, leaving IRTC in Grab mode will make a socket error.

Oct 31 05:24:20 lbtmu05 irc[32627]: GetFilter result status: ERROR [Error] IRSocket::GetFilter::res -> Could not read from socket.

BS9123 irtc.20081031.123051.fits

BS9122 irtc.20081031.123504.fits

BS9139 irtc.20081031.123805.fits

BS9126 irtc.20081031.124202.fits

Blank Sky 2.0 sec irtc.20081031.124306.fits

12:44 Cube of 20 x 2.0 sec blank sky frames 20081031_00117_cube.fit

12:45 Cube of 20 x 1.6 sec blank sky frames 20081031_00118_cube.fit

Blank Sky 1.6 sec irtc.20081031.124621.fits

BS9126 irtc.20081031.124731.fits

12:48 pointto BS9126 gs=0 TelM="ACTIVE" - Guider image is 0.6 arcsec FWHM.

A pair to measure the rotator center on IRTC:

BS9126 PA=0 irtc.20081031.125732.fits

BS9126 PA=180 irtc.20081031.125850.fits

Close

13:05 Close Dome, wind light, T=6.4 degC, D=-2.3 degC.

13:13 Moved all the images to /Repository/IRTC_Data/081031/ and cleaned up the ownership and permissions.

-- JohnHill - 31 Oct 2008
Topic revision: r4 - 31 Oct 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback